OpenStudy (anonymous):

Ok i understood it finally in regular problems but they had to go and give me a word problem and im confused now i thought i started it right but when i keep working through it i get no answers that match the possible ones

OpenStudy (anonymous):

Gina, Sam and Robby all rented movies from the same video store. They each rented some dramas, comedies and documentaries. Gina rented 11 movies total. Sam rented twice as many dramas, three times as many comedies, and twice as many documentaries as Gina. He rented 27 movies total. If Robby rented 19 movies total with the same number of dramas, twice as many comedies, and twice as many documentaries as Gina, how many movies of each type did Gina rent? 3 dramas, 5 comedies and 3 documentaries 2 dramas, 6 comedies and 3 documentaries 1 dramas, 4 comedies and 6 documentaries 4 dramas, 3 comedies and 4 documentaries

OpenStudy (anonymous):

I have to use substitution or elimnation

OpenStudy (anonymous):

i started out with these equations from it : a+b+c=11 2a+3b+2c=27 a+2b+2c=19 but, im not sure if these are right or not

OpenStudy (anonymous):

Or at least tell me if i started it off right with those equations ^^^^

OpenStudy (anonymous):

Let x represent the number of dramas Let y represent the number of comedies Let z represent the number of documentaries Gina: x + y + z = 11 Sam: 2x + 3y + 2z = 27 Robby: x + 2y +2z = 19

OpenStudy (anonymous):

a basically used a b and c but ok

OpenStudy (anonymous):

Rearrange the first equation for x: x = 11 - y - z

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

After rearranging you can put this equation in (2) or (3) and eliminate for y or z. Substitute x = 11 - y - z into (2) 2(11 - y - z) + 3y + 2z = 27 22 - 2y - 2z + 3y + 2z = 27 y = 27 - 22 y = 5 YAY that's the value of y :P Substitute y = 5 into (1) x + y + z = 11 x + 5 + z = 11 x + z = 6 x = 6 - z Substitute x = 6 - z and y = 5 into (3) x + 2y + 2z = 19 6 - z + 2y + 2z = 19 2y + z = 13 z = 13 - 2y z = 13 - 2(5) z = 3 There's your "z" value Substitute y = 5 and z = 3 into (1) to solve for x x + y + z = 11 x + 5 + 3 = 11 x = 11 - 8 x = 3 Check if your values satisfy the equation: LS = 2x + 3y + 2z = 2(3) + 3(5) + 2(3) = 6 + 15 + 6 = 27 RS = 27 Since LS = RS, the values of x, y and z satisfy the equation. Therefore, x = 3, y = 5 and z = 3. WELL?

OpenStudy (anonymous):

wow thank you so much

OpenStudy (anonymous):

ahah no problem. my back hurts now

OpenStudy (anonymous):

awe im sorry :(

OpenStudy (anonymous):

aha no worries.

OpenStudy (anonymous):

ok and that is what some good ole ibuprofen is for

OpenStudy (anonymous):

huh, yeah sure. :P